You can use the equation N = k log A to estimate the number of species N that live in a region of area A. The parameter k is determined by the conditions in the region. In a rain forest, 3200 species live in 600 km2
 How many species would remain if half of the forest area were destroyed by logging and farming?
a. 2853
b. 600
c. 3690

Answer by rothauserc(4718)

N = k * log(A)
:
3200 = k * log(600)
:
3200 = k * 2.7782
:
k = 3200 / 2.7782 = 1151.8249
:
N = 1151.8259 * log(300) = 2853
:
************
answer is a
************
: